The Lopez-Portillo Company has $10 million in assets, 80 percent financed by debt and 20 percent financed by common stock. The interest rate on the debt is 15 percent, and the stock book value is $10 per share. President Lopez-Portillo is considering two financing plans for an expansion to $15 million in assets. Under Plan A, the debt-to-total-assets ratio will be maintained, but new debt will cost 18 percent! New stock will be sold at $10 per share. Under Plan B, only new common stock at $10 per share will be issued. The tax rate is 40 percent.

a. If EBIT is 15 percent on total assets, compute earnings per share (EPS) before the expansion (current) and under the two alternatives after expansion. (Round the final answers to 2 decimal places.) EPS Current $ Plan A $ Plan B $

b. What is the DFL under each of the three plans? (Round the final answers to 2 decimal places.) DFL Current X Plan A X Plan B X

c. Using the formula provided in the chapter, calculate the EBIT/EPS indifference point between Plan A and Plan B. (Enter the answers in dollars not in millions.) EBIT $

d. If shares could be sold at $20 each due to increased expectations for the firm’s sales and earnings, what impact would this have on EPS for the two expansion alternatives? Compute EPS for each. (Round the final answers to 2 decimal places.) EPS Plan A Shares $ Plan B Shares $ e. Calculate the EBIT/EPS indifference point at the new share price. (Enter the answers in dollars not in millions.) EBIT $
Answer #1
Capital Structure ($ in Millions)
Particulars Current Plan Plan A Plan B
Total Assets 10 15 15
Old Debt 8 (0.8 * 10) 8 (0.8 * 10) 8 (0.8 * 10)
New Debt - 4 (0.8 * 5) -
Old Equity 2 (0.2 * 10) 2 (0.2 * 10) 2 (0.2 * 10)
New Equity - 1 (0.2 * 5) 5 (1 * 5)
Calculation of EBIT & EPS ($ in Millions)
Particulars Current Plan Plan A Plan B
EBIT (15% * Total Assets) 1.50 2.25 2.25
Less: Interest on old debt at 15% 1.20 1.20 1.20
Less: Interest on new debt at 18% - 0.72 -
EBT 0.30 0.33 1.05
Less: Taxes at 40% 0.12 0.132 0.42
EAT 0.18 0.198 0.63
Number of equity shares ($10 per share) 0.20 0.30 0.70

a. EPS = EAT / Number of equity shares

Current Plan = 0.18 / 0.20 = $0.90

Plan A = 0.198 / 0.30 = $0.66

Plan B = 0.63 / 0.70 = $0.90

b. DFL = EBIT / EBT

Current Plan = 1.50 / 0.30 = 5

Plan A = 2.25 / 0.33 = 6.82

Plan B = 2.25 / 1.05 = 2.14

c. EBIT/EPS indifference point = [(X - InterestA) (1 - Tax)] / Number of equity shares in Plan A = [(X - InterestB) (1 - Tax)] / Number of equity shares in Plan B

Where: X = EBIT indifference level

: [(X - 1.92) (1 - 0.40)] / 0.30 = [(X - 1.20) (1 - 0.40)] / 0.70

By solving above equation, we get X = $2,460,000

d. EPS if shares could be sold at $20:

Particulars Plan A Plan B
EAT $198,000 $630,000
Number of shares 250,000 450,000
EPS $0.792 $1.40

Plan A number of share = 200,000 shares(current) + (20% * 5,000,000) / $20 = 200,000 + 50,000 = 250,000

Plan B number of share = 200,000 shares(current) + 5,000,000 / $20 = 200,000 + 250,000 = 450,000

EBIT/EPS indifference point at the new share price :

[(X - 1.92) (1 - 0.40)] / 0.250 = [(X - 1.20) (1 - 0.40)] / 0.450

By solving above equation, we get X = $2,820,000